OpenStudy (anonymous):

Help! I give medals:) If a regulation soccer ball has a circumference of 70cm then what are the diameter radius surface area and volume.Show work

OpenStudy (anonymous):

@Jim766 @gaara438125 @ChrisNeedsHelp @thomaster do u think one of u could help me plezz :)

OpenStudy (jim766):

circumference = 2 pi r 70 = 2(3.14) r solve for r

geerky42 (geerky42):

C = 2πr So find radius from circumference then use the value of radius to find surface area \(SA = 4\pi r^2\), and volume \(V = \dfrac{4}{3}\pi r^3\)

OpenStudy (anonymous):

r=63.72??

geerky42 (geerky42):

To find r, divide both sides by 2π or 2(3.14)

OpenStudy (anonymous):

r=11.13

OpenStudy (anonymous):

or would u round n r would equal 12 @geerky42

OpenStudy (jim766):

go ahead and use that

OpenStudy (anonymous):

@Jim766 i get SA=576pi n v=2304pi is that correct?

OpenStudy (anonymous):

n diameter is 24?

OpenStudy (jim766):

diameter would be twice the radis

OpenStudy (jim766):

SA = 4(3.14)(11.13)^2= 1555.89

OpenStudy (anonymous):

i used 12